A balloon has a volume of 25.0L on a day when the temperature is 30.0 °C. If the temperature at night falls to 10.0 °C, what will be the volume of the balloon if the pressure remains constant?


1
Expert's answer
2022-04-07T13:12:39-0400

"PV=nRT \>\>by \>ideal\>gas\>law"

"P_{1}V_{1}=nRT_{1}\>\>equation (1)"

"P_{2}V_{2}=nRT_{2}\>\>equation (2)"

Pressure remains constant and also number of moles n and universal gas constant R is also constant.

"\frac{V_{1}}{T_{1}}=\frac{nR}{P}"

"\frac{V_{2}}{T_{2}}=\frac{nR}{P}"

"\frac{V_{1}}{T_{1}}=\frac{V_{2}}{T_{2}}"

"\frac{25L}{30^oC}=\frac{V_{2}}{10^oC}"

"V_{2}=8.33L"

"Volume\> of \>balloon \>is \>8.33 \>L"